# OWL2Vec4OA: Tailoring Knowledge Graph Embeddings for Ontology Alignment

### v1

@@ 1 Introduction
The ontology matching community has contributed to the evolution of ontology alignment systems for the last twenty years with the organization of the annual Ontology Alignment Evaluation Initiative (OAEI) [36, 35]. In recent years there has been a shift from traditional systems [11], using lexical and structural techniques, to systems using machine learning and large language models. Prominent examples include LogMap-ML [5], BertMap [21], DeepAlignment [29], VeeAlign [26], SORBETMatcher [15] and OLaLa [24]. The OAEI, with the new Bio-ML track [23, 22], has also evolved accordingly to attract and systematically evaluate such systems.
@@
In this paper, we present OWL2Vec4OA, an extension of the ontology embedding system OWL2Vec* [4] tailored to the ontology alignment task. OWL2Vec* projects a given ontology into a graph, randomly walks the graph to generate sequences of entities, and runs the language model Word2vec [32] to generate embeddings of both entity URIs and words. OWL2Vec4OA, unlike OWL2Vec*, relies on a potentially incomplete or inaccurate ontology alignments to bridge a given set of input ontologies. When projecting the ontologies and performing the random walks to create entity sequences, the confidence value of these seed mapping are used to bias the random walks (i.e., edges with higher confidence values will have higher chances to be visited). Hence, OWL2Vec4OA allows for a tighter connection of the input ontologies given a set of seed mappings, while giving preference to edges with higher confidence. [^1] Edges projected from ontology axioms are given the highest confidence. OWL2Vec4OA currently relies on the ontology matching systems LogMap [27] and AML [13] to produce seed mappings.
@@ 2 Preliminaries
are key for embedding systems like RDF2vec [37], OWL2Vec* and node2vec [16] to capture the structural and contextual information of a graph or network, so that the system can learn about the relationships between different nodes and their local neighborhoods. The PageRank algorithm [30] revolutionized web page ranking by employing a random walk model on the web’s hyperlink structure. It simulates a "random surfer" traversing a graph of web pages, computing page importance based on the probability of the surfer landing on each page. This approach effectively captures the web’s complex link topology to determine page significance, becoming a cornerstone in information retrieval and influencing various fields beyond web search. Random walks allow to process massive graphs without needing to consider all possible paths. Seminal works like DeepWalk [34] and node2vec[16] utilize random walks to generate node sequences for training embeddings, effectively capturing network topology. Wei [43] proposed an extension of the Metropolis-Hastings algorithm for sampling from large-scale networks, introducing strategies such as early rejection and biased sampling. RDF2vec [37] introduced random walk-based embeddings for RDF knowledge graphs, later extended by Steenwinckel [39] with new walk extraction strategies. Cochez et al. [7] also introduced biased walk strategies to RDF2vec with twelve different edge weighting functions.
@@ 3 Related Work
These diverse approaches demonstrate the complexity of ontology alignment and the variety of techniques employed [12]. Nonetheless, ontology matching tools can now process more efficiently even the most complex ontologies, including those with hundreds of thousands of classes, encompassing billions of possible connections. Recent advancements in ontology matching have incorporated machine learning, including embedding-based techniques and (large) language models, showing promising results in improving alignment accuracy (e.g., [5], [21]). These approaches leverage vector representations of ontological elements to capture semantic relationships more effectively than traditional methods. By utilizing language models or domain-specific embedding algorithms, these techniques can identify nuanced similarities between concepts, potentially leading to more accurate and comprehensive ontology alignments. This new generation of systems can broadly be categorized into three categories. (i) Direct embedding comparison: Methods like ERSOM [44] and DeepAlignment [29] calculate distances between concept embeddings directly. (ii) Supervised mapping classifiers: VeeAlign [26], MEDTO [18], LogMap-ML [5], and SORBET [15] train classifiers using concept embeddings as input. This approach adds a layer of learning specific to the OM task but still relies on independent embeddings for the input ontologies. (iii) Based on language models: Recent methods such as BERTMap [21], BERTSubs [3] and OLaLa [24] rely on language models to implement task-specific models. For instance, BERTMap fine-tunes pre-trained language models (PLMs) using ontology-specific information, while BERTSubs focuses on subsumption mapping prediction using context-based information.
@@ 4 Ontology embeddings with OWL2Vec4OA
Unlike our predecessor OWL2Vec*, OWL2Vec4OA builds a single graph taking as input the graph projections of the ontologies to be aligned and a set of seed mappings. Mappings (i.e., $\langle e,\allowbreak e^{\prime},\allowbreak r,\allowbreak c\rangle$) have a direct graph representation as triples, for example, in Figure 1 the following mapping ($r=$ equivalence) was identified $\langle$ HeLiS:Fructose, owl:equivalentClass, obo:FOODON_03301305 (fructose) $\rangle$ with confidence $c$ = 0.9. OWL2Vec4OA assigns a weight to each edge or link as follows: (i) $1.0$ if the edge was derived from ontology axioms; and (ii) $c$ if the edge was derived from a mapping. The output is a labeled weighted graph $G=(V,E)$, where $E$ is the set of edges built from the projected RDF triples and $V$ is the set of vertices composed by the subjects and objects in these triples. The function $weight(l)$ returns the weight for a given edge $l$, while the function $uri(e)$ returns the URI of a given entity.
Input: Weighted graph $G=(V,E)$; seed entities $S$; walk depth $wd$;
@@
             Append $uri(s)$ to $current\_walk$;
@@
                   Append $uri(l)$ to $current\_walk$; // URI of the link
                   Append $uri(v)$ to $current\_walk$; // URI of the vertex
@@
OWL2Vec4OA, inspired by Cochez et al. [7], implements the biased random walker summarised in Algorithm 1. The algorithm takes as input a weighted labeled graph $G$ and a set of seed entities $S$ and performs (biased) random walks of depth $wd$ starting from each of the seed entities. It optionally iterates over the seed entities more than once to allow for different walks for the same seed entity. In our setting, the seed entities represent the entities involved in the seed mappings computed in the alignment step. This way the walks are tailored to the alignment task, without the need of exploring the whole input ontologies. The bias in the random walk takes into account the weight assigned to each of the edges or links ($l$) in the graph to assign a probability to each of the potential paths. Given $E_{u}=\{l_{i}=(u,v_{i})\text{~{}with~{}}i=1..n\}$ and $l_{i}$ an outer edge for $u$, the probability for each edge is computed as in Equation 1.
$Pr(l_{j}=(u,v_{j}))=\frac{weight(l_{j})}{\sum_{i=1}^{n}weight(l_{i})}$ (1)
@@ 5 Evaluation
We have performed a preliminary evaluation of the suitability of the embeddings computed by OWL2Vec4OA in an ontology alignment tasks. Particularly we have used the datasets provided by the OAEI’s 2023 Bio-ML track [^3] Bio-ML Challenge [23]: https://krr-oxford.github.io/OAEI-Bio-ML/. Bio-ML 2023 Datasets: https://zenodo.org/records/8193375. The Bio-ML track included several tasks (e.g., OMIM-ORDO, NCIT-DOID, SNOMED-NCIT-Pharm and SNOMED-NCI-Neoplas), involving biomedical ontologies with tens of thousands of classes, and reference alignments based on Mondo [41] and UMLS [1].
@@
We have applied OWL2Vec4OA embeddings into the local matching tasks of Bio-ML. Mappings are scored and ranked according to the cosine similarity of the computed URI embeddings for the entities in the mapping. We have computed OWL2Vec4OA embeddings for different walk depths and iterations. We set the Word2Vec parameters epochs and embedding dimension to 70 and 100, respectively. The experiments were conducted on a High-Performance Computing cluster with access to up to 48 CPUs, using the Slurm workload manager to ensure efficient resource allocation and job scheduling Generated resources are available in Zenodo [40].
@@
Table 1 shows the results over the OMIM-ORDO for different sets of seed mappings as introduced in Section 4. We also used as seed the mappings provided as training and validation in Bio-ML, as shown in the first row, using training-validation mappings in isolation did not lead to promising results given their reduced size. The set of seed mappings leading to the best results was the union of LogMapout and AML mappings, with LogMapout $\cap$ AML and LogMapover leading to similar results. The impact of additional iterations over the seed entities did not lead to a significantly increased performance.
@@
Increasing the walk depths has a positive impact in the tasks NCIT-DOID and SNOMED-NCIT-Neoplas; while for OMIM-ORDO and SNOMED-NCIT-Pharm longer paths seem to add noise to the embeddings. This is inline with the results obtained in the original OWL2Vec* paper [4] where longer paths did not seem to lead to better results. It is worth mentioning that longer paths also increase the computation times. In the near future, we plan to perform an extended evaluation to better understand the impact of longer walks.
@@ Acknowledgements
Work funded by the Ministry of Science and Education of the Republic of Azerbaijan. This work has also been partially supported by the Academy of Medical Sciences Network Grant (Neurosymbolic AI for Medicine, NGR1\1857) and the project "XAI4SOC: Explainable Artificial Intelligence for Healthy Aging and Social Wellbeing" funded by the Agencia Estatal de Investigación (AEI), the Spanish Ministry of Science, Innovation and Universities and the European Social Funds (PID2021-123152OB-C22).

### v2

@@ 1 Introduction
The ontology matching community has contributed to the evolution of ontology alignment systems for the last twenty years with the organization of the annual Ontology Alignment Evaluation Initiative (OAEI) [36, 35]. In recent years there has been a shift from traditional systems [11], using lexical and structural techniques, to systems using machine learning and (large) language models. Prominent examples include LogMap-ML [5], BertMap [21], DeepAlignment [29], VeeAlign [26], SORBETMatcher [15] and OLaLa [24]. The OAEI, with the new Bio-ML track [23, 22], has also evolved accordingly to attract and systematically evaluate such systems.
@@
In this paper, we present OWL2Vec4OA, an extension of the ontology embedding system OWL2Vec* [4] tailored to the ontology alignment task. OWL2Vec* projects a given ontology into a graph, randomly walks over the graph to generate sequences of entities, and runs the language model Word2Vec [32] to generate embeddings of both entity URIs and words. OWL2Vec4OA, unlike OWL2Vec*, relies on potentially incomplete or inaccurate ontology alignments to bridge a given set of input ontologies. When projecting the ontologies and performing the random walks to create entity sequences, the confidence value of these seed mapping are used to bias the random walks (i.e., edges with higher confidence values will have higher chances to be visited). Hence, OWL2Vec4OA allows for a tighter connection of the input ontologies given a set of seed mappings, while giving preference to edges with higher confidence. [^1] Edges projected from ontology axioms are given the highest confidence. OWL2Vec4OA currently relies on the ontology matching systems LogMap [27] and AML [13] to produce seed mappings.
@@ 2 Preliminaries
are key for embedding systems like RDF2Vec [37], OWL2Vec* and node2vec [16] to capture the structural and contextual information of a graph or network, so that the system can learn about the relationships between different nodes and their local neighborhoods. The PageRank algorithm [30] revolutionized web page ranking by employing a random walk model on the web’s hyperlink structure. It simulates a “random surfer” traversing a graph of web pages, computing page importance based on the probability of the surfer landing on each page. This approach effectively captures the web’s complex link topology to determine page significance, becoming a cornerstone in information retrieval and influencing various fields beyond web search. Random walks allow to process massive graphs without needing to consider all possible paths. Seminal works like DeepWalk [34] and node2vec[16] utilize random walks to generate node sequences for training embeddings, effectively capturing network topology. Wei [43] proposed an extension of the Metropolis-Hastings algorithm for sampling from large-scale networks, introducing strategies such as early rejection and biased sampling. RDF2Vec [37] introduced random walk-based embeddings for RDF knowledge graphs, later extended by Steenwinckel [39] with new walk extraction strategies. Cochez et al. [7] also introduced biased walk strategies to RDF2Vec with twelve different edge weighting functions.
@@ 3 Related Work
These diverse approaches demonstrate the complexity of ontology alignment and the variety of techniques employed [12]. Nonetheless, ontology matching tools can now process more efficiently even the most complex ontologies, including those with hundreds of thousands of classes, encompassing billions of possible connections. Recent advancements in ontology matching have incorporated machine learning, including embedding-based techniques and (large) language models, showing promising results in improving alignment accuracy (e.g., [5], [21]). These approaches leverage vector representations of ontological elements to capture semantic relationships more effectively than traditional methods. By utilizing language models or domain-specific embedding algorithms, these techniques can identify nuanced similarities between concepts, potentially leading to more accurate and comprehensive ontology alignments. This new generation of systems can broadly be categorized into three categories. (i) Direct embedding comparison: Methods like ERSOM [44] and DeepAlignment [29] calculate distances between concept embeddings directly. (ii) Supervised mapping classifiers: VeeAlign [26], MEDTO [18], LogMap-ML [5], and SORBET [15] train classifiers using concept embeddings as input. This approach adds a layer of learning specific to the OM task but still relies on independent embeddings for the input ontologies. (iii) Based on language models: Recent methods such as BERTMap [21], BERTSubs [3] and OLaLa [24] rely on language models to implement task-specific models. For instance, BERTMap fine-tunes pre-trained language models (PLMs) using synonyms from the ontologies, while BERTSubs focuses on subsumption mapping prediction using context-based information that is transformed into text by templates.
@@ 4 Ontology embeddings with OWL2Vec4OA
Unlike our predecessor OWL2Vec*, OWL2Vec4OA builds a single graph taking as input the graph projections of the ontologies to be aligned and a set of seed mappings. Mappings (i.e., $\langle e,\allowbreak e^{\prime},\allowbreak r,\allowbreak c\rangle$) have a direct graph representation as triples, for example, in Figure 1 the following mapping ($r=$ equivalence) was identified $\langle$ HeLiS:Fructose, owl:equivalentClass, obo:FOODON_03301305 (fructose) $\rangle$ with confidence $c$ = 0.9. OWL2Vec4OA assigns a weight to each edge or link as follows: (i) $1.0$ if the edge was derived from ontology axioms; and (ii) $c$ if the edge was derived from a mapping. The output is a labeled weighted graph $G=(V,E,U,W)$, where $E$ is the set of edges built from the projected RDF triples, $V$ is the set of vertices composed by the subjects and objects in these triples, $U$ is the the set of URIs associated to the vertices and edges, and $W$ is the set of weights associated to the edges. The function $weight(G,l)$ returns the weight for a given edge $l$, while the function $uri(G,e)$ returns the URI of a given entity $e$.
Input: Weighted graph $G=(V,E,U,W)$; seed entities $S$; walk depth $wd$;
@@
             Append $uri(G,e)$ to $current\_walk$;
@@
                   Append $uri(G,l)$ to $current\_walk$; // URI of the link
                   Append $uri(G,v)$ to $current\_walk$; // URI of the vertex
@@
OWL2Vec4OA, inspired by Cochez et al. [7], implements the biased random walker summarised in Algorithm 1. The algorithm takes as input a weighted labeled graph $G$ and a set of seed entities $S$ and performs (biased) random walks of depth $wd$ starting from each of the seed entities. It optionally iterates over the seed entities more than once to allow for different walks for the same seed entity. In our setting, the seed entities represent the entities involved in the seed mappings computed in the alignment step. This way the walks are tailored to the alignment task, without the need of exploring the whole input ontologies. The bias in the random walk takes into account the weight assigned to each of the edges or links ($l$) in the graph $G$ to assign a probability to each of the potential paths. Given $E_{u}=\{l_{i}=(u,v_{i})\text{~{}with~{}}i=1..n\}$ and $l_{i}$ an outer edge for $u$, the probability for each edge is computed as in Equation 1.
$Pr(l_{j}=(u,v_{j}))=\frac{weight(G,l_{j})}{\sum_{i=1}^{n}weight(G,l_{i})}$ (1)
@@ 5 Evaluation
We have performed a preliminary evaluation of the suitability of the embeddings computed by OWL2Vec4OA in ontology alignment tasks. Particularly we have used the datasets provided by the OAEI’s 2023 Bio-ML track [^3] Bio-ML Challenge [23]: https://krr-oxford.github.io/OAEI-Bio-ML/. Bio-ML 2023 Datasets: https://doi.org/10.5281/zenodo.8193375. The Bio-ML track included several tasks (e.g., OMIM-ORDO, NCIT-DOID, SNOMED-NCIT-Pharm and SNOMED-NCI-Neoplas), involving biomedical ontologies with tens of thousands of classes, and reference alignments based on Mondo [41] and UMLS [1].
@@
We have applied OWL2Vec4OA embeddings into the local matching tasks of Bio-ML. Mappings are scored and ranked according to the cosine similarity of the computed URI embeddings for the entities in the mapping. We have computed OWL2Vec4OA embeddings for different walk depths and iterations. We fixed the Word2Vec hyperparameters — the number of epochs and embedding dimension to 70 and 100, respectively. The experiments were conducted on a High-Performance Computing cluster with access to up to 48 CPUs, using the Slurm workload manager to ensure efficient resource allocation and job scheduling. Generated resources are available in Zenodo [40].
@@
Table 1 shows the results over the OMIM-ORDO for different sets of seed mappings as introduced in Section 4. We also used as seed the mappings provided as training and validation in Bio-ML, as shown in the first row, using training-validation mappings in isolation did not lead to promising results given their reduced size. The set of seed mappings leading to the best results was the union of LogMapout and AML mappings, with LogMapout $\cap$ AML and LogMapover leading to similar results. The results also show that OWL2Vec4OA is also able to handle noisy set of seed mappings like LogMapover. The impact of additional iterations over the seed entities did not lead to a significantly increased performance.
@@
Increasing the walk depths has a positive impact in the tasks NCIT-DOID and SNOMED-NCIT-Neoplas; while for OMIM-ORDO and SNOMED-NCIT-Pharm longer paths seem to add noise to the embeddings. This is inline with the results obtained in the original OWL2Vec* paper [4] where longer paths did not seem to lead to better results. It is worth mentioning that longer paths also increase the computation times. In the near future, we plan to perform an extended evaluation to better understand the impact of longer walks on different ontologies and matching tasks.
@@ Acknowledgements
This research is funded by the Ministry of Education and Science of Azerbaijan Republic with support from City St George’s, University of London. This work has also been partially supported by the Academy of Medical Sciences Network Grant (Neurosymbolic AI for Medicine, NGR1\1857), the project "XAI4SOC: Explainable Artificial Intelligence for Healthy Aging and Social Wellbeing" funded by the Agencia Estatal de Investigación (AEI), the Spanish Ministry of Science, Innovation and Universities and the European Social Funds (PID2021-123152OB-C22), the EPSRC project OntoEm (EP/Y017706/1), and the EU Projects: RE4DY (101058384, HORIZON-CL4-2021), Plooto (101092008, HORIZON-CL4-2022), SM4RTENANCE (101123490, DIGITAL-2022), and Tec4MaasEs (101138517, HORIZON-CL4-2023).